OpenStudy (idealist10):

Solve for y: 2y-y^2=(x+c)^2.

OpenStudy (cwrw238):

Hint: complete the square on LHS

OpenStudy (idealist10):

I'm not too good at completing the square. So I got -y^2+2y=-(y^2-2y), now what?

OpenStudy (cwrw238):

ok - (y^2 - 2y) = (x + c)^2 y^2 - 2y = -(x = c)^2 (y -1 )^2 - 1 = -(x + c)^2 (y - 1)^2 = 1 - (x + c)^2 can you continue from here?

OpenStudy (idealist10):

Let me work it out.

OpenStudy (idealist10):

Thanks!

OpenStudy (cwrw238):

yw

OpenStudy (idealist10):

So y-1=sqrt(1-(x+c)^2), right?

OpenStudy (cwrw238):

right - 0h 2 solutions - positive and negative

OpenStudy (cwrw238):

+/- what you've written

OpenStudy (idealist10):

:)

OpenStudy (cwrw238):

so add 1 to both sides to get y and you're done

OpenStudy (idealist10):

Wait, so if I add 1 to the right side, I got y=1+/-sqrt(1-(x+c)^2), right?

OpenStudy (idealist10):

@cwrw238